# Gamma family characterization and an alternative proof of Gini estimator unbiasedness

### v1

@@ Abstract
In this paper, we derive a general representation for the expectation of the (upward adjusted) Gini coefficient estimator in terms of the Laplace transform of the underlying distribution. This representation leads to a characterization of the gamma family within the class of nonnegative scale families, based on a stability property of exponentially tilted distributions. As an application, we provide an alternative proof of the unbiasedness of the Gini coefficient estimator under gamma populations.
Keywords. Gamma distribution, Gini coefficeint, Gini estimator, unbiased estimator.
@@ 1 Introduction
A well-known issue is the finite-sample bias of the classical Gini estimator. To address it, Deltas, (2003) proposed an upward-adjusted estimator that corrects small-sample bias under particular distributional assumptions. Its unbiasedness has since been established for several distribution families, most notably the gamma family; see Baydil et al., (2025) and Vila and Saulo, 2025a ; Vila and Saulo, 2025b .
This paper makes two complementary contributions. First, we derive a general expression for the expectation of the Gini estimator for any nonnegative, absolutely continuous distribution with finite mean. This approach naturally connects to the Gini mean difference (GMD) studied in Vila et al., (2024), highlighting how the distributional properties of the underlying population influence the Gini-based measures of variability and inequality. Second, we show that a natural scale-family invariance condition on the associated exponentially tilted distribution characterizes the gamma family.
This characterization yields a unified proof of the unbiasedness of the Gini estimator under gamma populations and clarifies the fundamental role of scale invariance in explaining the prominence of the gamma family in Gini estimation.
The remainder of the paper is organized as follows. In Section 2 we review basic properties and representations of the Gini coefficient. In Section 3 we derive a general expression for the expectation of the Gini estimator. In Section 4 we establish a characterization of the gamma family within the class of scale families. In Section 5 we apply these results to obtain an alternative proof of the unbiasedness of the Gini estimator under gamma populations. Finally, in Section 6, we provide some concluding remarks.
@@ 2 The Gini coefficient
Let $X_{1},X_{2}$ be independent and identically distributed (i.i.d.) non-negative random variables with common distribution as a nonnegative, absolutely continuous random variable $X$, and finite mean $\mu=\mathbb{E}[X]>0$. The Gini coefficient (Gini,, 1936) is defined as
@@
The Gini coefficient $G$ of a nonnegative, absolutely continuous random variable $X$ can be characterized as
$G={\int_{0}^{\infty}F_{X}(x)[1-F_{X}(x)]{\rm d}x\over\mu},$
where $\mu=\mathbb{E}[X]>0$ is the mean of $X$, and $F_{X}$ is the CDF of $X$.
@@
$|x-y|=\int_{0}^{\infty}\left[1-\mathds{1}_{\{x\leqslant t,y\leqslant t\}}-\mathds{1}_{\{x\geqslant t,y\geqslant t\}}\right]{\rm d}t.$ (2.1)
3 A simple expression for the expectation of the Gini estimator
In this section (Theorem 3.1), we derive a general expression for the expectation of the (upward adjusted) Gini coefficient estimator $\widehat{G}$, initially proposed by Deltas, (2003).
Let $X_{1},\ldots,X_{n}$ be i.i.d. copies of $X$. The Gini coefficient estimator of $G$ is defined by
$\widehat{G}=\dfrac{\displaystyle{1\over\binom{n}{2}}\sum_{1\leqslant i<j\leqslant n}|X_{i}-X_{j}|}{2\overline{X}},$
where $\overline{X}=\sum_{k=1}^{n}X_{k}/n$ denotes the sample mean.
@@ 3 A simple expression for the expectation of the Gini estimator
Let $X_{1},\ldots,X_{n}$ be independent copies of a nonnegative, absolutely continuous random variable $X$ with finite mean $\mu=\mathbb{E}[X]>0$ and CDF $F_{X}$. Then
@@
where $\mathscr{L}_{X}(z)=\mathbb{E}\left[\exp\{-zX\}\right]$ denotes the Laplace transform of $X$. For each $z>0$, $G(Y_{z})$ is the Gini coefficient of the random variable $Y_{z}$, whose CDF is given by
$F_{Y_{z}}(t)=\frac{\mathbb{E}\left[\exp\{-zX\}\mathds{1}_{\{X\leqslant t\}}\right]}{\mathscr{L}_{X}(z)},\quad t>0.$ (3.1)
@@
$={n\over 2}\int_{0}^{\infty}\mathbb{E}\left[|X_{1}-X_{2}|\exp\{-z(X_{1}+X_{2})\}\right]\mathscr{L}_{F}^{n-2}(z){\rm d}z,$
@@
$\mathbb{E}\left[\widehat{G}\right]={n\over 2}\int_{0}^{\infty}\left\{\int_{0}^{\infty}\left\{\mathscr{L}_{F}^{2}(z)-\mathbb{E}^{2}\left[\exp\{-zX\}\mathds{1}_{\{X\leqslant t\}}\right]-\mathbb{E}^{2}\left[\exp\{-zX\}\mathds{1}_{\{X\geqslant t\}}\right]\right\}{\rm d}t\,\mathscr{L}_{F}^{n-2}(z)\right\}{\rm d}z.$
@@
$\mathbb{E}\left[\widehat{G}\right]={n}\int_{0}^{\infty}\left\{\int_{0}^{\infty}F_{Y_{z}}(t)\left[1-F_{Y_{z}}(t)\right]{\rm d}t\,\mathscr{L}_{X}^{n}(z)\right\}{\rm d}z.$ (3.2)
@@
$\int_{0}^{\infty}F_{Y_{z}}(t)\left[1-F_{Y_{z}}(t)\right]{\rm d}t=\mathbb{E}[Y_{z}]G(Y_{z}),$ (3.3)
@@ 4 Characterization of the gamma family
5 Unbiasedness of the Gini estimator
@@
Hence, $E[Y_{z}]=\alpha/(\lambda+z)$, $\mathscr{L}_{X}(z)=\lambda^{\alpha}/(\lambda+z)^{\alpha}$, and since the Gini coefficient is scale invariant, $G(Y_{z})=G(X)=G$. Applying Theorem 3.1, we obtain

### v2

# On the bias of the Gini estimator: Poisson and geometric cases, a characterization of the gamma family, and unbiasedness under gamma distributions

@@ Abstract
In this paper, we derive a general representation for the expectation of the Gini coefficient estimator in terms of the Laplace transform of the underlying distribution, together with the mean and the Gini coefficient of its exponentially tilted version. This representation leads to a new characterization of the gamma family within the class of nonnegative scale families, based on a stability property under exponential tilting. As direct applications, we show that the Gini estimator is biased for both Poisson and geometric populations and provide an alternative, unified proof of its unbiasedness for gamma populations. By using the derived bias expressions, we propose plug-in bias-corrected estimators and assess their finite-sample performance through a Monte Carlo study, which demonstrates substantial improvements over the original estimator. Compared with existing approaches, our framework highlights the fundamental role of scale invariance and exponential tilting, rather than distribution-specific algebraic calculations, and complements recent results in Baydil et al., (2025) and Vila and Saulo, 2025a ; Vila and Saulo, 2025b .
Keywords. Gamma distribution, Gini coefficient, Gini estimator, unbiased estimator.
@@ 1 Introduction
A well-known issue is the finite-sample bias of the classical Gini estimator. To address it, Deltas, (2003) proposed an upward-adjusted estimator that corrects small-sample bias under particular distributional assumptions. Its unbiasedness has since been established for several distribution families, most notably the gamma family; see Baydil et al., (2025) and Vila and Saulo, 2025a ; Vila and Saulo, 2025b . These results raise a natural question: what structural properties of the underlying distribution explain the unbiasedness of the Gini estimator, and to what extent does this phenomenon extend beyond the gamma family?
This paper addresses these questions by developing a general representation for the expectation of the Gini estimator that applies to any non-degenerate nonnegative distribution with finite mean. The proposed representation expresses the expectation of the estimator in terms of the Laplace transform of the underlying distribution and the Gini coefficient of its exponentially tilted version. This formulation naturally connects with the Gini mean difference framework studied in Vila et al., (2024) and provides a transparent way to analyze bias properties across different distributional families.
Within this framework, we obtain four main contributions. First, we show explicitly that the Gini estimator is biased for Poisson populations, deriving an exact integral representation for its expectation and sharp bounds for the resulting bias. Second, we extend this analysis to geometric populations, where we again establish bias and provide closed-form and hypergeometric representations for the expectation of the estimator. These results demonstrate that bias persists even for simple, classical discrete distributions. Third, we show that a scale-invariance property of exponentially tilted distributions characterizes the gamma family, yielding a unified and conceptually simple proof of the unbiasedness of the Gini estimator under gamma populations. Fourth, we use the derived bias expressions of the Gini estimators (Poisson and geometric) to propose bias-corrected versions through a plug-in approach. A Monte Carlo study is carried out to assess the finite-sample behavior of both the original and the corrected estimators. The simulation results indicates that the bias-corrected estimators exhibit a markedly improved performance.
The remainder of the paper is organized as follows. In Section 2, we review basic properties and representations of the Gini coefficient. In Section 3, we derive a general expression for the expectation of the Gini estimator. In Sections 3.1 and 3.2, we apply this result to Poisson and geometric populations, respectively, establishing finite-sample bias in both cases. In Section 4, we establish a characterization of the gamma family based on exponential tilting, and Section 4.1 uses this characterization to obtain an alternative proof of the unbiasedness of the Gini estimator for gamma populations. In Section 5, we propose bias-corrected estimators for the Poisson and geometric cases and report results from a simulation study assessing the finite-sample performance of all estimators. Finally, in Section 6, we conclude with a discussion of implications and directions for future research.
@@ 2 The Gini coefficient
Let $X_{1},X_{2}$ be independent and identically distributed (i.i.d.) random variables with common distribution as a non-degenerate nonnegative random variable $X$, and finite mean $\mu=\mathbb{E}[X]>0$. The Gini coefficient (Gini,, 1936) is defined as
@@
The Gini coefficient $G$ of a nonnegative random variable $X$ can be characterized as
$G={\int_{0}^{\infty}F_{X}(x^{-})[1-F_{X}(x^{-})]{\rm d}x\over\mu},$
where $\mu=\mathbb{E}[X]>0$ is the mean of $X$, $F_{X}$ is the CDF of $X$, and $F_{X}(x^{-})=\mathbb{P}(X<x)$.
@@
$|x-y|=\int_{0}^{\infty}\left[\mathds{1}_{(x,\infty)}(t)+\mathds{1}_{(y,\infty)}(t)-2\mathds{1}_{(x,\infty)}(t)\mathds{1}_{(y,\infty)}(t)\right]{\rm d}t,$ (2.1)
where $\mathds{1}_{A}(t)$ is the indicator function of a set $A$. ∎
Proposition 2.2.
The Gini coefficient $G$ of a discrete random variable $X\in\{0,1,\ldots\}$ can be characterized as
$G={\sum_{k=0}^{\infty}F_{X}(k)[1-F_{X}(k)]\over\mu}.$
Proof.
Since
$F_{X}(x^{-})=F_{X}(\lfloor x\rfloor)=F_{X}(k),\quad x\in[k,k+1),$
where $\lfloor x\rfloor$ denotes the floor function, the function $F_{X}(x^{-})$ is constant on each interval $[k,k+1)$. Consequently,
$\int_{0}^{\infty}F_{X}(x^{-})\,[1-F_{X}(x^{-})]\,\mathrm{d}x=\sum_{k=0}^{\infty}\int_{k}^{k+1}F_{X}(\lfloor x\rfloor)\,[1-F_{X}(\lfloor x\rfloor)]\,\mathrm{d}x.$
Therefore, by Proposition 2.1, the proof follows. ∎
3 A general expression for the expectation of the Gini estimator
In this section (Theorem 3.2), we derive a general expression for the expectation of the Gini coefficient estimator $\widehat{G}$.
Let $X_{1},\ldots,X_{n}$ be i.i.d. copies of $X$. Following De la Peña et al., (2025), we define an estimator of the Gini coefficient $G$ by
$\widehat{G}=\begin{cases}\dfrac{{1\over\binom{n}{2}}\sum_{1\leqslant i<j\leqslant n}|X_{i}-X_{j}|}{2\overline{X}},&\text{if}\ \sum_{k=1}^{n}X_{k}>0\ \text{a.s.},\\[8.5359pt] 0,&\text{if}\ \sum_{k=1}^{n}X_{k}=0\ \text{a.s.},\end{cases}$ (3.1)
where $\overline{X}=\sum_{k=1}^{n}X_{k}/n$ denotes the sample mean. Note that, in the case where $\sum_{k=1}^{n}X_{k}>0$ almost surely, $\widehat{G}$ coincides with the upward adjusted Gini coefficient estimator originally proposed by Deltas, (2003).
Remark 3.1.
In an earlier version of this work, the following result was established only for i.i.d. absolutely continuous random variables, independently of the results obtained in De la Peña et al., (2025). Becoming aware of that reference motivated us to extend this result to i.i.d. random variables in a more general setting, which can be further extended, following the approach of De la Peña et al., (2025), by removing the assumption that the variables involved are identically distributed.
@@ 3 A general expression for the expectation of the Gini estimator
Let $X_{1},\ldots,X_{n}$ be independent copies of a nonnegative random variable $X$ with finite mean $\mu=\mathbb{E}[X]>0$ and CDF $F_{X}$. Then
@@
where $\mathscr{L}_{X}(z)=\mathbb{E}\left[\exp\{-zX\}\right]$ denotes the Laplace transform of $X$. For each $z>0$, $G(Y_{z})$ is the Gini coefficient of an exponentially tilted (or Esscher-transformed) random variable $Y_{z}$ (Butler,, 2007), whose CDF is given by
$F_{Y_{z}}(t)=\frac{\mathbb{E}\left[\exp\{-zX\}\mathds{1}_{(0,t]}(X)\right]}{\mathscr{L}_{X}(z)},\quad t>0.$ (3.2)
@@
$={n\over 2}\int_{0}^{\infty}\mathbb{E}\left[|X_{1}-X_{2}|\exp\{-z(X_{1}+X_{2})\}\right]\mathscr{L}_{X}^{n-2}(z){\rm d}z,$
@@
$\mathbb{E}\left[\widehat{G}\right]={n}\int_{0}^{\infty}\left\{\int_{0}^{\infty}\mathbb{E}\left[\exp\{-zX\}\mathds{1}_{(0,t)}(X)\right]\left\{\mathscr{L}_{X}(z)-\mathbb{E}\left[\exp\{-zX\}\mathds{1}_{(0,t)}(X)\right]\right\}{\rm d}t\,\mathscr{L}_{X}^{n-2}(z)\right\}{\rm d}z.$
@@
$\mathbb{E}\left[\widehat{G}\right]={n}\int_{0}^{\infty}\left\{\int_{0}^{\infty}F_{Y_{z}}(t^{-})\left[1-F_{Y_{z}}(t^{-})\right]{\rm d}t\,\mathscr{L}_{X}^{n}(z)\right\}{\rm d}z.$ (3.3)
@@
$\int_{0}^{\infty}F_{Y_{z}}(t^{-})\left[1-F_{Y_{z}}(t^{-})\right]{\rm d}t=\mathbb{E}[Y_{z}]G(Y_{z}),$ (3.4)
@@
Remark 3.3.
Observe that $\mathbb{E}[Y_{z}]$ and $G(Y_{z})$ in Theorem 3.2 satisfy the following relations
$\mathbb{E}[Y_{z}]={\mathbb{E}[X\exp\{-zX\}]\over\mathscr{L}_{X}(z)}=-{\partial\log\left(\mathscr{L}_{X}(z)\right)\over\partial z}$
and
$G(Y_{z})={\mathbb{E}\left[|X_{1}-X_{2}|\exp\{-z(X_{1}+X_{2})\}\right]\over 2\mathbb{E}\left[X\exp\{-z(X)\}\right]\mathscr{L}_{X}(z)}=A_{1}(z)A_{2}(z)\,G,$
where $G=G(X)$ is the Gini coefficient of $X$, $A_{1}$ is a weighted difference factor,
$A_{1}(z)\equiv{\mathbb{E}\left[|X_{1}-X_{2}|\exp\{-z(X_{1}+X_{2})\}\right]\over\mathbb{E}|X_{1}-X_{2}|}\leqslant 1,$
because $\exp\{-z(X_{1}+X_{2})\}$ is bounded above by 1, and $A_{2}$ is a mean adjustment factor,
$A_{2}(z)\equiv{\mu\over\mathbb{E}[X\exp\{-zX\}]\mathscr{L}_{X}(z)}\geqslant 1,$
since, by the generalized Chebyshev sum inequality, $X$ and $\exp\{-zX\}$ are negatively correlated.
By applying Theorem 3.2 together with Remark 3.3, we have
Corollary 3.4.
Under the assumptions of Theorem 3.2,
$\mathbb{E}\left[\widehat{G}\right]=\left[-n\int_{0}^{\infty}{\partial\log\left(\mathscr{L}_{X}(z)\right)\over\partial z}\,A_{1}(z)A_{2}(z)\mathscr{L}_{X}^{\,n}(z)\,\mathrm{d}z\right]G,$
where $G=G(X)$ is the Gini coefficient of $X$.
3.1 Biased Gini estimator in Poisson populations
If $X\sim\mathrm{Poisson}(\lambda)$, $\lambda>0$, then the exponentially tilted variable $Y_{z}$ is Poisson with mean $\lambda\exp\{-z\}$. By Proposition 2.2, the Gini coefficient of $X$ is
$G(X)=\exp\{-2z\}\big[I_{0}(2\lambda)+I_{1}(2\lambda)\big],$
where $I_{0}$ and $I_{1}$ denote the modified Bessel functions of the first kind. By the same argument,
$G(Y_{z})=\exp\{-2\lambda\exp\{-z\}\}\big[I_{0}(2\lambda\exp\{-z\})+I_{1}(2\lambda\exp\{-z\})\big].$
Moreover, since $\mathbb{E}[Y_{z}]=\lambda\exp\{-z\}$ and $\mathscr{L}_{X}(z)=\exp\{\lambda(\exp\{-z\}-1)\}$, the change of variables $w=\exp\{-z\}$ in Theorem 3.2 yields
$\mathbb{E}\left[\widehat{G}\right]=n\lambda\exp\{-nz\}\int_{0}^{1}\exp\{(n-2)\lambda w\}\big[I_{0}(2\lambda w)+I_{1}(2\lambda w)\big]\,\mathrm{d}w.$
Although this integral has no closed form in terms of elementary or classical special functions, the integrand is continuous on $[0,1]$; hence the integral is finite and $\mathbb{E}\left[\widehat{G}\right$] is well defined.
Moreover, since $1\leqslant I_{0}(2\lambda w)+I_{1}(2\lambda w)\leqslant I_{0}(2\lambda)+I_{1}(2\lambda)$, we get
${\exp\{-2\lambda\}-\exp\{-n\lambda\}\over(1-{2\over n})}\leqslant\mathbb{E}\left[\widehat{G}\right]\leqslant{1-\exp\{-(n-2)\lambda\}\over(1-{2\over n})}\,G,\quad n\geqslant 1,n\neq 2,$
and since $\int_{0}^{1}\bigl[I_{0}(2\lambda w)+I_{1}(2\lambda w)\bigr]\,\mathrm{d}w=[{I_{0}(2\lambda)+I_{1}(2\lambda)-1}]/({2\lambda}),$ we have
$\mathbb{E}\left[\widehat{G}\right]=G-\exp\{-2\lambda\},\quad n=2.$
In other words, the estimator $\widehat{G}$ is biased for Poisson populations, with bias
$\text{Bias}(\widehat{G},G)=n\lambda\exp\{-n\lambda\}\int_{0}^{1}\exp\{(n-2)\lambda w\}[I_{0}(2\lambda w)+I_{1}(2\lambda w)]{\rm d}w-\exp\{-2\lambda\}[I_{0}(2\lambda)+I_{1}(2\lambda)]$ (3.5)
satisfying
${\exp\{-2\lambda\}-\exp\{-n\lambda\}\over(1-{2\over n})}-G\leqslant\text{Bias}(\widehat{G},G)\leqslant\left[{1-\exp\{-(n-2)\lambda\}\over(1-{2\over n})}-1\right]G,\quad n\geqslant 1,n\neq 2,$
and
$\text{Bias}(\widehat{G},G)=-\exp\{-2\lambda\},\quad n=2.$
3.2 Biased Gini estimator in geometric populations
If $X\sim\mathrm{Geometric}(p)$, $p\in(0,1)$, on $\{0,1,\ldots\}$, then the exponentially tilted variable $Y_{z}$ is geometric with parameter $1-(1-p)\exp\{-z\}$. By Proposition 2.2,
$G(X)=\frac{1}{2-p},\quad G(Y_{z})=\frac{1}{1+(1-p)\exp\{-z\}}.$
Moreover, since
$\mathbb{E}[Y_{z}]=\frac{(1-p)e^{-z}}{1-(1-p)\exp\{-z\}}\quad\text{and}\quad\mathscr{L}_{X}(z)=\frac{p}{1-(1-p)\exp\{-z\}},$
the change of variables $w=(1-p)\exp\{-z\}$ yields the representation of $\mathbb{E}[\widehat{G}$] in Theorem 3.2 as
$\mathbb{E}\left[\widehat{G}\right]=np^{n}\int_{0}^{1-p}{1\over(1-w)^{n+1}(1+w)}\,\mathrm{d}w={1\over 2}\left[{}_{2}F_{1}\left(1,n;n+1;{p\over 2}\right)-p^{n}\,_{2}F_{1}\left(1,n;n+1;{1\over 2}\right)\right],$
where in the last equality we have used the standard integral representation of the Gauss hypergeometric function (see Chapter 15, Item 15.3. in Abramowitz and Stegun,, 1972):
$_{2}F_{1}(a,b;c;z)=\frac{\Gamma(c)}{\Gamma(b)\Gamma(c-b)}\int_{0}^{1}t^{b-1}(1-t)^{c-b-1}(1-zt)^{-a}\,{\rm d}t,\quad c>b>0.$
Since, for $0\leqslant w\leqslant 1-p$, it holds that $G\leqslant 1/(1+w)\leqslant 1$, it follows that
$(1-p^{n})\,G=\left[np^{n}\int_{0}^{1-p}{1\over(1-w)^{n+1}}\,\mathrm{d}w\right]G\leqslant\mathbb{E}\left[\widehat{G}\right]\leqslant np^{n}\int_{0}^{1-p}{1\over(1-w)^{n+1}}\,\mathrm{d}w={1-p^{n}}.$
In other words, the estimator $\widehat{G}$ is biased for Geometric populations, with bias given by
$\text{Bias}(\widehat{G},G)={1\over 2}\left[{}_{2}F_{1}\left(1,n;n+1;{p\over 2}\right)-p^{n}\,_{2}F_{1}\left(1,n;n+1;{1\over 2}\right)\right]-{1\over 2-p}$ (3.6)
which satisfies
$-p^{n}G\leqslant\text{Bias}(\widehat{G},G)\leqslant 1-p^{n}-G.$
@@ 4 Characterization of the gamma family
4.1 Unbiased Gini estimator in gamma populations
@@
Hence, $E[Y_{z}]=\alpha/(\lambda+z)$, $\mathscr{L}_{X}(z)=\lambda^{\alpha}/(\lambda+z)^{\alpha}$, and since the Gini coefficient is scale invariant,
$G(Y_{z})=G(X)=G=\frac{\Gamma(2\alpha+1)}{2^{2\alpha}\Gamma(\alpha+1)^{2}},$
where the last identity is well known in the literature and can be easily obtained from Proposition 2.1. Applying Theorem 3.2, we obtain
@@
5 Simulation study
This section reports a Monte Carlo study assessing the finite-sample performance of the Gini estimator in (3.1) (uncorrected) and its bias-corrected version obtained by subtracting an estimate of the finite-sample bias. The correction is implemented in a plug-in fashion using maximum likelihood (ML) estimation of the distributional parameter.
Let $X_{1},\ldots,X_{n}$ be an i.i.d. sample from a Poisson or Geometric distribution. The uncorrected estimator is given in (3.1), that is,
$\widehat{G}=\begin{cases}\dfrac{{1\over\binom{n}{2}}\sum_{1\leqslant i<j\leqslant n}|X_{i}-X_{j}|}{2\overline{X}},&\text{if}\ \sum_{k=1}^{n}X_{k}>0\ \text{a.s.},\\[8.5359pt] 0,&\text{if}\ \sum_{k=1}^{n}X_{k}=0\ \text{a.s.},\end{cases}$
where $\overline{X}=\sum_{k=1}^{n}X_{k}/n$ denotes the sample mean. The corrected estimator is defined as
[…]
For the Poisson model, we consider $X\sim\mathrm{Poisson}(\lambda)$ with $\lambda\in\{0.5,1,2,5,10\}$. For each Monte Carlo replication and each pair $(n,\lambda)$, the ML estimator of $\lambda$ is $\widehat{\lambda}=\bar{X}$. The plug-in bias correction is computed using the analytical expression in Eq. (3.5).
For the geometric model, we generate $X\sim\mathrm{Geometric}(p)$ on $\{0,1,\ldots\}$ with $p\in\{0.1,0.2,0.4,0.6,0.8\}$. For each replication and each pair $(n,p)$, the ML estimator is $\widehat{p}=1/(1+\bar{X})$ and the bias correction is obtained from Eq. (3.6).
Sample sizes are set to $n\in\{25,50,75,100\}$, and each configuration is replicated $R$ times, with $R$ sufficiently large to ensure stable Monte Carlo estimates (in our implementation, R=10, [^0] R=10{,}000). For each replication $r$, we compute $(\widehat{G}_{r},\widehat{G}_{c,r})$.
Performance is summarized using the relative bias and the root mean squared error (RMSE), defined respectively as
$\mathrm{RelBias}(\widehat{G}_{e})=\frac{1}{R}\sum_{r=1}^{R}\frac{\widehat{G}_{e,r}-G}{G},\qquad\mathrm{RMSE}(\widehat{G}_{e})=\left(\frac{1}{R}\sum_{r=1}^{R}(\widehat{G}_{e,r}-G)^{2}\right)^{1/2},$
where $\widehat{G}_{e}$ denotes either $\widehat{G}$ or $\widehat{G}_{c}$ and $G$ is the true Gini coefficient under the data-generating model.
Figure 1 reports the Monte Carlo results for the Poisson model. Panels (a) and (b) display the relative bias and RMSE as functions of the sample size $n$, for fixed values of $\lambda$, while panels (c) and (d) show the same measures as functions of $\lambda$, for fixed sample sizes. From this figure, we observe that the bias-corrected estimator displays reduced relative bias across all configurations. Regarding efficiency, the RMSE decreases with increasing $n$ for both estimators, with quite similar performances.
(a) Relative bias vs $n$ (Poisson)
(b) RMSE vs $n$ (Poisson)
(c) Relative bias vs $\lambda$ (Poisson)
(d) RMSE vs $\lambda$ (Poisson)
Figure 1: Monte Carlo results for the Poisson model: relative bias and RMSE for the uncorrected estimator (Gb, Eq. (3.1)) and the bias-corrected version (Gcorr).
Figure 2 presents the corresponding results for the geometric model. Panels (a) and (b) show relative bias and RMSE as functions of $n$, whereas panels (c) and (d) depict these measures as functions of the success probability $p$. As in the Poisson case, the bias-corrected estimator displays reduced relative bias across all configurations, with similar RMSE performances.
(a) Relative bias vs $n$ (Geometric)
(b) RMSE vs $n$ (Geometric)
(c) Relative bias vs $p$ (Geometric)
(d) RMSE vs $p$ (Geometric)
Figure 2: Monte Carlo results for the geometric model: relative bias and RMSE for the uncorrected estimator (Gb, Eq. (3.1)) and the bias-corrected version (Gcorr).